Circular No. 7609

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V4334 SAGITTARII K. Hinkle and R. Joyce, Kitt Peak National Observatory (KPNO),

report infrared photometry of V4334 Sgr (cf. IAUC 6322) obtained on two different telescopes and instruments on Mar. 13.6 UT: “Observers Hinkle and Joyce at the KPNO 4-m telescope (using the Phoenix spectrograph in imaging mode) report [1.083 μm] = 13.6, [1.205 μm] = 13.0, [1.647 μm] = 9.4, [2.230 μm] = 6.1, and [3.226 μm] = 3.3. Observers M. Brown and G. Tiede at the KPNO 2.1-m telescope (using SQIID) report J = 12.60, H = 9.36, and K = 6.42. This confirms the reported brightening (cf. IAUC 7592). The infrared photometry, except for [1.083 μm], is consistent with a 680-K blackbody that is unchanged over the last year (cf. IAUC 7422).”

T. R. Geballe, Gemini Observatory; B. Smalley, V. H. Tyne, and A. Evans, Keele University; and S. P. S. Eyres, University of Central Lancashire, obtained medium-resolution J and K-band spectra of V4334 Sgr on Mar. 21.6 UT at the U.K. Infrared Telescope: “We estimate J = 13.2 and K = 6.7 from the spectra, confirming the brightening in the J band observed by Lederle and Kimeswenger (IAUC 7592), but possibly indicating some fading since their measurement. The brightening at J is due to a stronger continuum; the flux in the He line at 1.083 μm has decreased markedly since our previous observation in Aug. 2000. We urge continued monitoring of this object.”

1998 SF_{36}

R. P. Binzel, Massachusetts Institute of Technology, reports obtaining 0.8- to 2.5-μm spectroscopic measurements of 1998 SF_{36} on Feb. 19.5 and Mar. 28.6 UT using the NASA Infrared Telescope Facility 3-m reflector on Mauna Kea: “Combined with Kitt Peak 4-m visible spectrum measurements (cf. IAUC 7598), complete coverage is available over 0.5–2.5 μm. These data fully reveal the features of an S-type minor-planet spectrum whose 1-μm-feature center is at 0.99 ± 0.01 μm and 2- to 1-μm ‘band area ratio’ is 0.40 ± 0.02, placing it within the S(IV) category of Gaffey et al. (1993, Icarus 106, 573). The minor-planet reflectance spectrum resembles that of laboratory measurements of LL chondrite meteorites (Gaffey 1976, J. Geophys. Res. 81, 905; all spectra normalized to unity at 0.55 μm), although it appears more steeply red-sloped by an average reflectance of 0.20 per μm. 1998 SF_{36} appears to have the surface composition of a reddened ordinary chondrite meteorite.”

2001 April 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green